LostLegacy06:

Can someone attempt to explain this: x/(x+3)+(x+2)/(x+5)

minustempo:

Since the denominators are different, you need to first find a common denominator

LostLegacy06:

LostLegacy06:

Ok-

minustempo:

Here the least common denominator is just (x+3)(x+5)

LostLegacy06:

wouldn't i need to multiply that to the top?

minustempo:

you would need to multiply something on the top to get the denominator to (x+3)(x+5), yes

LostLegacy06:

ultimate confusion

minustempo:

so you know that a/b + c/d is not equal to (a+c)/(b+d) right?

LostLegacy06:

mhm

minustempo:

It's equal to (ad + bc)/(bd)

minustempo:

because if you want the common denominator as bd, then a/b becomes ad/bd, and c/d becomes bc/bd Now both denominators are equal, so you can add ad/bd + bc/bd = (ad+bc)/bd

minustempo:

so x/(x+3) + (x+2)/(x+5) = (x(x+5) + (x+2)(x+3)/((x+3)(x+5))

LostLegacy06:

I somewhat understand...but im still confused-

LostLegacy06:

so the x(x+5) is by itself?

minustempo:

Sorry I don't know latex It's x(x+5) + (x+2)(x+3)


(x+3)(x+5)

LostLegacy06:

ohhhh ok that looks and makes more sense

LostLegacy06:

makeshift fraction sign did wonders

minustempo:

lol

LostLegacy06:

so I can cross out the (x+5) and (x+3) from the top and bottom right?

minustempo:

You don't cross out, you divide the top and bottom by the same number. but you can't do that here because the x(x+5) term doesn't have an (x+3) and the (x+2)(x+3) term doesn't have an (x+5) term

LostLegacy06:

yeah I've hit rock bottom

minustempo:

lol

LostLegacy06:

i'm going to try to figure out how to do this based on what you said and if I do I will let you know, gracias <3

minustempo:

I don't think you can simplify any farther though, but sure

LostLegacy06:

well yeah I still have to figure out how you got there xD

minustempo:

Sure lmao